4) The SI unit for mass is ___ and for length is 5) a) 1.345 s = hr b) 247.5 kg pg
Answer #1

ans)

from above data that

4)

we know that the

the SI unit for mass is kilogram symbol (kg)

the SI unit for length is meter (m)

5)

a)1.345 s=......

here we convert to seconds per hour

we know that

1 sec=0.000278 hour

so we get

1.345 sec=0.0003739 hour

b)

247.5 kg=........pg

here convert kilograms to pico grams

1 kilogram=1*10^15

so we get

247.5 kilograms=1*10^15*247.5

=2.47*10^17 picograms